# ScholarMate: A Mixed-Initiative Tool for Qualitative Knowledge Work and Information Sensemaking

### v1

@@ Abstract.
Synthesizing knowledge from large document collections is a critical yet increasingly complex aspect of qualitative research and knowledge work. While AI offers automation potential, effectively integrating it into human-centric sensemaking workflows remains challenging. We present ScholarMate, an interactive system designed to augment qualitative analysis by unifying AI assistance with human oversight. ScholarMate enables researchers to dynamically arrange and interact with text snippets on a non-linear canvas, leveraging AI for theme suggestions, multi-level summarization, and contextual naming, while ensuring transparency through traceability to source documents. Initial pilot studies indicated that users value this mixed-initiative approach, finding the balance between AI suggestions and direct manipulation crucial for maintaining interpretability and trust. We further demonstrate the system’s capability through a case study analyzing 24 papers. By balancing automation with human control, ScholarMate enhances efficiency and supports interpretability, offering a valuable approach for productive human-AI collaboration in demanding sensemaking tasks common in knowledge work.
1. Introduction & Motivation
@@
Researchers deal with large amounts of text‐based data in their daily workflows, such as reviewing academic publications and conducting qualitative data analysis on interview transcripts. Making sense of these large corpora presents escalating challenges due to the sheer volume of text data being collected and produced. Traditional approaches, such as manual open coding and thematic labeling, require extensive human effort and time, making rapid adjustments to emerging patterns that are challenging when faced with large corpora. This demanding manual process relies heavily on the researcher’s interpretive skill to discern patterns and relationships (Tashman and Edwards, 2011), limiting the scale and efficiency of analysis possible (Lim, ).
@@
Recent advances in large language models (LLMs) have demonstrated their capacity to summarize, cluster, and infer patterns from large textual datasets (Brown et al., 2020). Despite these capabilities, purely automated solutions can struggle with context awareness and interpretive flexibility—qualities that are critical in qualitative research domains (Huang et al., 2025; Hung et al., 2023). Moreover, integrating AI raises concerns about automation bias, the potential for diminished researcher cognitive engagement (Ye et al., 2025), and the risk of over-reliance on outputs that may lack nuance or contain errors (Hou et al., 2023; Huang et al., 2025). Three core challenges persist in large‐scale AI-assisted analysis: (1) Maintaining transparency and traceability when automated summarization systems are introduced (Kryscinski et al., 2020), (2) ensuring researchers remain actively engaged (Hou et al., 2023) rather than passively accepting outputs, and (3) supporting creative insight through flexible exploration of emergent connections during the research process (Anderson et al., 2024).
@@
We implement this in ScholarMate, an LLM-assisted system designed to support researchers conducting qualitative research work and facilitating the text-based sensemaking process. ScholarMate allows researchers to import a text corpus and manipulate snippets on a non-linear canvas interface, enabling them to build working theories of related documents. This visual re-arrangement process specifically supports human-AI collaborative sensemaking. Furthermore, ScholarMate explores how HCI can shape future work practices involving AI partners, aiming for meaningful and productive outcomes in complex knowledge work.
@@ 3. Design Goals
DG2: Ensure Interpretability and Transparency. AI models often act as opaque “black boxes”, undermining researcher confidence and constraining critical assessment (Ye et al., 2025). To address this, ScholarMate explicitly connects each AI-generated suggestion, whether proposing to assign evidence to an existing theme or create a new one, with supporting evidence nodes directly on the interactive canvas (Amershi et al., 2019). A visual preview shows the outcome before the user commits. Transparent explanations, supported by this preview and direct traceability, help researchers understand how and why suggestions are made, facilitating informed acceptance or revision (Kulesza et al., 2015) and promoting active cognitive engagement.
DG3: Promote Ethical Integration and Maintain Trust. AI tools can reinforce biases or generate plausible but incorrect information (“hallucinations”), introducing ethical challenges and undermining research validity (Ye et al., 2025; Holstein et al., 2019). Therefore, ScholarMate facilitates critical assessment by providing clear reasoning for its suggestions and ensuring direct traceability to the source evidence. By transparently communicating its limitations and supporting necessary human oversight, the system maintains ethical standards and builds user trust.
@@ 4. Probe Design
We developed a web-based prototype, known as a technology probe (Hutchinson et al., 2003), to demonstrate the initial functionalities of ScholarMate. The prototype is currently deployed and can be accessed online333https://scholar-mate-eta.vercel.app/. ScholarMate is built using NextJS, with react-pdf-viewer and ReactFlow components, and designed to support interactive, mixed-initiative qualitative analysis through a flexible, non-linear canvas-based interface. It addresses challenges researchers face in systematically analyzing data while balancing automated assistance with human analytic control. Grounded in mixed-initiative interaction principles (Horvitz, 1999) and established qualitative research methodologies (Braun and Clarke, 2006), ScholarMate ensures automated suggestions remain transparent and subject to human validation, recognizing the critical role of human judgment in resolving ambiguities and refining interpretations.
@@ 4.1. Interactive Canvas
To navigate potentially complex canvas layouts while preserving relative spatial positioning of data, ScholarMate incorporates Semantic Zoom. Grounded in visualization principles like Shneiderman’s mantra – “overview first, zoom and filter, then details on demand” (Shneiderman, 2003) – for exploring large datasets, this allows fluid transitions between high-level thematic overviews and detailed evidence views. As a researcher zooms out, node text transforms into progressively shorter, LLM-generated summaries of original content. Conversely, zooming in reveals longer summaries or original text excerpts.
This feature combination directly supports our key design goals: blending user control with AI summaries achieves Mixed-Initiative Collaboration (DG1); clear visualization and adaptive summaries ensure Interpretability and Transparency (DG2); and empowering researchers with direct organizational control promotes Ethical AI Integration and Trust (DG3).
@@ 4.2. Working Canvas and Codebook View
ScholarMate provides two distinct but interconnected interfaces: the Working Canvas and the Codebook View. Both interfaces utilize ReactFlow technology, semantic zoom, node-based interactions, and maintain direct evidence-to-PDF highlighting capabilities.
The Working Canvas is designed to support active data analysis, enabling users to freely organize, edit, and refine evidence and theme nodes. It supports full interactivity, allowing detailed modifications, annotations, and iterative analysis, empowering users to co-create and dynamically explore emerging thematic insights. Conversely, the Codebook View offers a stable, read-only visualization of current themes and evidence. To facilitate comprehension without overwhelming users, the Codebook View limits the display to two pieces of evidence per theme. This interface emphasizes high-level theme exploration, providing a quick reference for users to check their understanding or synthesize existing themes.
This dual-interface approach supports Mixed-Initiative Collaboration (DG1), allowing both dynamic, AI-supported exploration and human-led refinement (Working Canvas), as well as ensuring transparency and interpretability of finalized thematic decisions (DG2). It further supports maintaining trust (DG3) by explicitly distinguishing exploratory work from the reporting stage, thus enhancing accountability.
@@ 4.3. AI-Assisted Features
Theme Suggestions: After a user selects a text excerpt, SCHOLARMATE analyzes it and the existing workspace themes. It then suggests either assigning the excerpt to a relevant existing theme or using it to establish a new one. Alternatively, users can choose to manually incorporate the selected excerpt directly onto the workspace, allowing them full control over its placement and connections within their analysis structure. AI suggestions are visibly distinguished by special UI indicators and transparently explained, supporting informed choices and ensuring researchers maintain ultimate control over their thematic decisions, whether leveraging AI assistance or proceeding manually.
@@
Contextual Theme Naming: Researchers can utilize AI-generated theme naming functionality that dynamically suggests theme names based on the aggregated context of connected evidence nodes. This feature supports the iterative nature of qualitative research by allowing researchers to continually refine theme names as their understanding of the data evolves. It helps researchers quickly capture and communicate the essence of thematic clusters, enhancing conceptual coherence and facilitating iterative sense-making.
@@
The AI-assisted features directly address all three design goals: facilitating mixed-initiative collaboration (DG1) by providing contextually relevant recommendations while maintaining researcher autonomy; ensuring interpretability and transparency (DG2) by clearly linking AI suggestions to source evidence; and promoting ethical AI integration and trust (DG3) through explicit communication of uncertainties and confidence levels.
@@ 5.1. Pilot Study
We conducted a preliminary evaluation of ScholarMate through a small pilot study with two graduate students (P1, P2), both with backgrounds in Human-Computer Interaction (HCI) and Data Visualization. We mainly aim to understand: (1) participants’ strategies for using ScholarMate in a qualitative research task and (2) their perceptions of its usefulness. Participants were asked to complete a thematic analysis task on two familiar research papers (Wang et al., 2025; Gorniak et al., 2024), pre-loaded into ScholarMate with researcher-provided highlights matching two predefined themes. Participants, however, were not informed of these pre-identified themes and were asked to independently identify themes using ScholarMate. Afterward, we administered NASA-TLX (Hart and Staveland, 1988) to measure perceived cognitive load and UMUX-LITE (Lewis et al., 2013) to assess system usability, followed by brief semi-structured interviews for qualitative feedback.
Quantitative Results. While our sample size was small, participants rated ScholarMate highly, with an average satisfaction score of 4.5 out of 5. The NASA-TLX scores indicated low-moderate cognitive demands (aggregated load: 2.41/7), primarily driven by low frustration (average 1.5/7), physical demand (average 1/7), and high perceived task performance (average 6/7). UMUX-LITE results showed strong usability ratings, indicating high participant agreement that the system met their requirements and was easy to use.
Qualitative Results. Participants identified both strengths and areas for improvement:
Beneficial AI Integration. Both participants appreciated AI-generated themes and insights, stating they streamlined the analysis and reduced coding effort: “If this can consistently assign the right theme, it saves a lot of time” (P2). AI suggestions were particularly helpful as editable initial points: “It’s useful to…use an AI-suggested theme and edit later” (P2). P1 found AI particularly effective at handling specific text excerpts.
Flexible Interaction. Participants liked the flexible, canvas-based interface, which accommodated diverse organizational styles (hierarchical for P1, flatter structures for P2). Overall, they found it highly compatible with their analysis workflows: “It seems like a way I would actually want to do qualitative research” (P2).
Limitations and Feature Requests. Participants suggested improving usability with better multi-node management and undo/history features. Concerns were noted about realistic scalability: “the scale… is different for an actual project” (P2). Regarding AI capabilities, P1 highlighted current limitations in the AI’s understanding of hierarchical relationships. Advanced suggestions included chatbot-like interactions and “top-down” AI guidance through user-specified analysis objectives at the session outset.
@@ 5.2. Use Case: Analyzing Vector Database Papers with ScholarMate
Notably, ScholarMate is not limited to matching predefined categories. Instead, its primary value is enabling iterative sensemaking and thematic exploration. The interactive workspace and semantic zoom capabilities empowered researchers with fluid transitions, moving smoothly from macro-level thematic overviews to micro-level analysis of individual excerpts., highlighting ScholarMate’s strength in supporting sophisticated analysis.
@@ 6. Limitations and Future Work
As ScholarMate incorporates interactions with LLMs, it faces inherent risks of AI error. To reduce uncritical reliance on AI, we designed ScholarMate to foster cognitive engagement by providing direct links to source excerpts, explanations for suggested themes, and easy-to-use editing capabilities. Future work will add visual differentiation of AI/human content for traceability and accountability. We also plan broader empirical studies with diverse users (e.g., researchers, analysts) to assess interactions, scalability, and workplace needs. To further streamline large-scale analysis, future work could explore integrating more automated tagging features, leveraging LLMs guided by user-defined codebooks, an approach shown to be methodologically feasible (Choksi et al., 2024). Enhanced privacy controls for sensitive user data are also planned.
@@ 7. Conclusion
We introduced ScholarMate, an LLM-assisted visual tool enhancing qualitative analysis through a mixed initiative canvas. Guided by design goals of human-AI collaboration (DG1), interpretability (DG2), and trust (DG3), ScholarMate aims to foster effective human-AI partnerships that augment, rather than automate, the nuanced process of qualitative research. Our initial case study and pilot evaluations suggest that ScholarMate can serve as a useful tool for textual data exploration and organization, successfully supporting sensemaking while maintaining the researcher’s critical cognitive engagement.
@@ Acknowledgements.
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2024-04348 and RGPIN-2024-06005].References
@@
- Choksi et al. (2024) Madiha Zahrah Choksi, Marianne Aubin Le Quéré, Travis Lloyd, Ruojia Tao, James Grimmelmann, and Mor Naaman. 2024. Under the (neighbor)hood: Hyperlocal Surveillance on Nextdoor. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 771, 22 pages. doi:10.1145/3613904.3641967
@@
Appendix A Appendix
A.1. Full Qualitative Results for ScholarMate
A.1.1. Beneficial AI Integration
A.1.2. Freeform Interaction
A.1.3. Limitation and Future Improvments

### v2

@@ Abstract.
Synthesizing knowledge from large document collections is a critical yet increasingly complex aspect of qualitative research and knowledge work. While AI offers automation potential, effectively integrating it into human-centric sensemaking workflows remains challenging. We present ScholarMate, an interactive system designed to augment qualitative analysis by unifying AI assistance with human oversight. ScholarMate enables researchers to dynamically arrange and interact with text snippets on a non-linear canvas, leveraging AI for theme suggestions, multi-level summarization, and evidence-based theme naming, while ensuring transparency through traceability to source documents. Initial pilot studies indicated that users value this mixed-initiative approach, finding the balance between AI suggestions and direct manipulation crucial for maintaining interpretability and trust. We further demonstrate the system’s capability through a case study analyzing 24 papers. By balancing automation with human control, ScholarMate enhances efficiency and supports interpretability, offering a valuable approach for productive human-AI collaboration in demanding sensemaking tasks common in knowledge work.
1. Introduction
@@
Researchers deal with large amounts of text‐based data in their daily workflows, such as reviewing academic publications and conducting qualitative data analysis on interview transcripts. Making sense of these large corpora presents escalating challenges due to the sheer volume of text data being collected and produced. Traditional approaches, such as manual open coding and thematic labeling, require extensive human effort and time, making rapid adjustments to emerging patterns that are challenging when faced with large corpora (Braun and Clarke, 2006). This demanding manual process relies heavily on the researcher’s interpretive skill to discern patterns and relationships (Tashman and Edwards, 2011), limiting the scale and efficiency of analysis possible (Lim, ).
@@
Recent advances in large language models (LLMs) have demonstrated their capacity to summarize, cluster, and infer patterns from large textual datasets (Brown et al., 2020). Despite these capabilities, purely automated solutions can struggle with context awareness and interpretive flexibility, qualities that are critical in qualitative research domains (Huang et al., 2025; Hung et al., 2023). Moreover, integrating AI raises concerns about automation bias, the potential for diminished researcher cognitive engagement (SKITKA et al., 1999), and the risk of over-reliance on outputs that may lack nuance or contain errors (Hou et al., 2023; Huang et al., 2025). Three core challenges persist in large‐scale AI-assisted analysis: (1) Maintaining transparency and traceability when automated summarization systems are introduced (Kryscinski et al., 2020), (2) ensuring researchers remain actively engaged (Hou et al., 2023) rather than passively accepting outputs, and (3) supporting creative insight through flexible exploration of emergent connections during the research process (Anderson et al., 2024).
@@
We implement this in ScholarMate, an LLM-assisted system designed to support researchers conducting qualitative research work and facilitating the text-based sensemaking process. ScholarMate allows researchers to import a text corpus and manipulate snippets on a non-linear canvas interface, enabling them to build working theories of related documents. This visual re-arrangement process specifically supports human-AI collaborative sensemaking. Furthermore, ScholarMate explores how human-computer interaction (HCI) can shape future work practices involving AI partners, aiming for meaningful and productive outcomes in complex knowledge work.
@@ 3. Design Goals
DG2: Ensure Interpretability of AI Reasoning. AI models often act as opaque ”black boxes,” which can hinder researchers’ ability to critically assess their outputs and reasoning processes (Miller, 2018; Ye et al., 2025). To illuminate the AI’s reasoning process, ScholarMate provides transparency into the basis of each suggestion. It explicitly links AI-generated suggestions to the specific evidence snippets that prompted them, offering a direct rationale. A visual preview further clarifies the immediate impact of accepting a suggestion. Making this rationale transparent allows researchers to critically evaluate the logic behind a suggestion, facilitating informed interaction: acceptance, revision, or rejection, based on their own understanding (Kulesza et al., 2015). This aligns with human-AI interaction guidelines promoting system intelligibility (Amershi et al., 2019).
DG3: Promote Output Validation and Ethical Use. AI tools can introduce errors through biases or hallucinations, posing ethical challenges and threatening research validity (Ye et al., 2025; Holstein et al., 2019). Therefore, ScholarMate incorporates features focused on validating AI-generated content and ensuring ethical integration. Direct traceability allows researchers to easily verify AI claims or summaries against original source documents, confirming accuracy. Transparently communicating AI limitations and mandating human oversight are crucial mechanisms to mitigate risks associated with potential errors or biases. This emphasis on verification and human control helps maintain ethical standards and fosters justified trust in the final research outcomes.
@@ 4. Probe Design
ScholarMate’s design was informed by observing qualitative researchers’ existing practices. We analyzed how they used informal spatial arrangements and annotations (often sketches) to cluster document excerpts and build understanding during their real-world workflows. Insights from these observations, combined with iterative internal design reviews focused on addressing identified researcher needs, guided the development of successive low- and high-fidelity prototypes.
This process led to the current web-based technology probe (Hutchinson et al., 2003), to demonstrate the initial functionalities of ScholarMate. The ScholarMate probe can be accessed online333https://scholar-mate-eta.vercel.app/. ScholarMate is built using NextJS, with react-pdf-viewer and ReactFlow components, and designed to support interactive, mixed-initiative qualitative analysis through a flexible, non-linear canvas-based interface. It addresses challenges researchers face in systematically analyzing data while balancing automated assistance with human analytic control. Grounded in mixed-initiative interaction principles (Horvitz, 1999) and established qualitative research methodologies (Braun and Clarke, 2006), ScholarMate ensures automated suggestions remain transparent and subject to human validation, recognizing the critical role of human judgment in resolving ambiguities and refining interpretations.
@@ 4.1. Interactive Canvas
To navigate potentially complex canvas layouts while preserving relative spatial positioning of data, ScholarMate incorporates Semantic Zoom. Grounded in visualization principles like Shneiderman’s mantra – “overview first, zoom and filter, then details on demand” (Shneiderman, 2003) – for exploring large datasets, this allows fluid transitions between high-level thematic overviews and detailed evidence views. As a researcher zooms out, evidence text transforms into progressively shorter, LLM-generated summaries of original content. Conversely, zooming in reveals longer summaries or original text excerpts. While evidence nodes adapt their content detail, theme nodes primarily scale visually with the zoom level, retaining their fixed text to ensure stable thematic anchors.
This combination of direct manipulation and AI-driven semantic zoom primarily supports Mixed-Initiative Collaboration (DG1) by blending user control with AI assistance. Furthermore, the crucial ability to trace any evidence snippet on the canvas directly back to its context in the original source document promotes Output Validation (DG3), allowing researchers to easily verify information and trust the evidence underpinning their analysis.
@@ 4.2. Working Canvas and Codebook View
ScholarMate provides two distinct but interconnected interfaces: the Working Canvas and the Codebook View. Both interfaces utilize ReactFlow component, semantic zoom, node-based interactions, and maintain direct evidence-to-PDF highlighting capabilities.
The Working Canvas is designed to support active data analysis, enabling users to freely organize, edit, and refine evidence and theme nodes. It supports full interactivity, allowing detailed modifications, annotations, and iterative analysis, empowering users to co-create with AI, and dynamically explore emerging thematic insights. Conversely, the Codebook View offers a stable, read-only visualization of current themes and evidence. To facilitate comprehension without overwhelming users, the Codebook View limits the display to two pieces of evidence per theme. This interface emphasizes high-level theme exploration, providing a quick reference for users to check their understanding or synthesize existing themes.
This dual-interface design directly enhances Mixed-Initiative Collaboration (DG1) by providing distinct spaces for dynamic, AI-supported exploration versus human-led refinement and finalization. Separating exploratory work (Working Canvas) from the stable, read-only synthesis (Codebook View) also promotes Ethical Use (DG3) by improving the accountability and trustworthiness of the final thematic structure.
@@ 4.3. AI-Assisted Features
Theme Suggestions: After a user selects a text excerpt, ScholarMate analyzes it and the existing workspace themes. It then suggests either assigning the excerpt to a relevant existing theme or using it to establish a new one. Alternatively, users can choose to manually incorporate the selected excerpt directly onto the workspace, allowing them full control over its placement and connections within their analysis structure. AI suggestions are visibly distinguished by special UI indicators and transparently explained, supporting informed choices and ensuring researchers maintain overall control over their thematic decisions, whether leveraging AI assistance or proceeding manually.
@@
Contextual Theme Naming: Researchers can utilize AI-generated theme naming functionality that suggests theme names based on the aggregated context of connected evidence nodes. This feature supports the iterative nature of qualitative research by allowing researchers to continually refine theme names as their understanding of the data evolves. It helps researchers quickly capture and communicate the essence of thematic clusters, enhancing conceptual coherence and facilitating iterative sense-making.
@@
These AI-assisted features embody Mixed-Initiative Collaboration (DG1) by offering suggestions while preserving user autonomy. Features like Theme Suggestions and Evidence-Grounded Theme Descriptions directly address Interpretability of AI Reasoning (DG2) by linking suggestions and keywords to the underlying evidence, explaining why they were proposed. Concurrently, the reliance on source evidence for suggestions and summaries necessitates and supports Output Validation (DG3), encouraging researchers to verify AI contributions against original documents.
@@ 5.1. Pilot Study
We conducted a preliminary evaluation of ScholarMate through a small pilot study with two graduate students (P1, P2), both with backgrounds in human-computer interaction and data visualization. We mainly aim to understand: (1) participants’ strategies for using ScholarMate in a qualitative research task and (2) their perceptions of its usefulness. Participants were asked to complete a thematic analysis task on two familiar research papers (Wang et al., 2025; Gorniak et al., 2024), pre-loaded into ScholarMate with researcher-provided highlights matching two predefined themes. Participants, however, were not informed of these pre-identified themes and were asked to independently identify themes using ScholarMate. Afterward, we administered NASA-TLX (Hart and Staveland, 1988) to measure perceived cognitive load and UMUX-LITE (Lewis et al., 2013) to assess system usability, followed by brief semi-structured interviews for qualitative feedback.
5.2. Quantitative Results
While our sample size was small, participants rated ScholarMate highly, with an average satisfaction score of 4.5 out of 5. The NASA-TLX scores indicated low-moderate cognitive demands (aggregated load: 2.41/7), primarily driven by low frustration (average 1.5/7), physical demand (average 1/7), and high perceived task performance (average 6/7). UMUX-LITE results showed strong usability ratings, indicating high participant agreement that the system met their requirements and was easy to use.
5.3. Qualitative Results
Participants identified both strengths and areas for improvement:
5.3.1. Beneficial AI Integration
5.3.2. Freeform Interaction
5.3.3. Limitation and Future Improvments
@@ 5.4. Use Case: Analyzing Vector Database Papers with ScholarMate
Notably, ScholarMate is not limited to matching predefined categories. Instead, its primary value is enabling iterative sensemaking and thematic exploration. The interactive workspace and semantic zoom capabilities empowered researchers with fluid transitions, moving smoothly from macro-level thematic overviews to micro-level analysis of individual excerpts, highlighting ScholarMate’s strength in supporting sophisticated analysis.
@@ 6. Limitations and Future Work
As ScholarMate incorporates interactions with LLMs, it faces inherent risks of AI error. To reduce over-reliance on AI, we designed ScholarMate to foster cognitive engagement by providing direct links to source excerpts, explanations for suggested themes, and easy-to-use editing capabilities. Future iterations could enhance the analytical workflow by supporting more complex, realistic hierarchical structures (e.g., themes, sub-themes, nested concepts) to better capture nuanced relationships within the data. Building upon existing interpretability features, we plan to investigate various interaction designs for human-AI collaboration, evaluating how they influence appropriate trust and reliance based on recent insights into user interactions with LLMs (Bo et al., 2024). Additionally, while initial feedback is promising, larger-scale user studies involving diverse researchers and analysts are essential to rigorously evaluate ScholarMate’s interactions, scalability, and impacts on real-world practices. Finally, to further streamline analysis, future work could integrate more automated tagging features guided by user-defined codebooks and implement enhanced privacy controls for sensitive data.
@@ 7. Conclusion
We introduced ScholarMate, an LLM-assisted visual interactive tool enhancing qualitative analysis through a mixed initiative canvas. Guided by design goals of supporting mixed-initiative human-AI collaboration (DG1), ensuring interpretability of AI reasoning (DG2), and promoting output validation and ethical use (DG3), ScholarMate aims to foster effective human-AI partnerships that augment, rather than automate, the nuanced process of qualitative research. Our initial case study and pilot evaluations suggest that ScholarMate can serve as a useful tool for textual data exploration and organization, successfully supporting sensemaking while maintaining the researcher’s critical cognitive engagement.
@@ Acknowledgements.
We would like to thank Prof. Michael Liut for supporting this work. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2024-04348 and RGPIN-2024-06005].References
@@
- Bo et al. (2024) Jessica Y Bo, Sophia Wan, and Ashton Anderson. 2024. To Rely or Not to Rely? Evaluating Interventions for Appropriate Reliance on Large Language Models. arXiv preprint arXiv:2412.15584 (2024).
@@
- Miller (2018) Tim Miller. 2018. Explanation in Artificial Intelligence: Insights from the Social Sciences. arXiv:1706.07269 [cs.AI] https://arxiv.org/abs/1706.07269
@@
- SKITKA et al. (1999) LINDA J. SKITKA, KATHLEEN L. MOSIER, and MARK BURDICK. 1999. Does automation bias decision-making? International Journal of Human-Computer Studies 51, 5 (1999), 991–1006. doi:10.1006/ijhc.1999.0252